two cubes of mass 3 g and 6 g with side 2 cm and 3 cm are kept on a table respectively. calculate the pressure experienced by the table


1
Expert's answer
2020-08-10T19:54:36-0400
"P=\frac{mg}{a^2}+\frac{Mg}{b^2}\\P=\frac{(0.003)(9.8)}{0.02^2}+\frac{(0.006)(9.8)}{0.03^2}\\P=139\ Pa"